OpenStudy (anonymous):

y varies inversely with x. If y=40 when x=16, find x when y=-5.

OpenStudy (anonymous):

Inverse variation is: \(y = \dfrac{k}{x}\) Plug in what we know: \(40 = \dfrac{k}{16}\) Now multiply 16 to both sides, what's 40 * 16?

OpenStudy (anonymous):

Can you multiply 40 * 16? @INarcolepsy

OpenStudy (anonymous):

640

OpenStudy (anonymous):

Yes, so the constant is 640. Plug in -5 for 'y' and 640 for 'k': \(y = \dfrac{k}{x}\) \(-5 = \dfrac{640}{x}\) Multiply 'x' to both sides; \(-5x = 640\) Divide -5 to both sides, what's 640 / -5?

OpenStudy (anonymous):

128

OpenStudy (anonymous):

Yes, that's our answer @INarcolepsy